# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 6.4.2. Pretraining Datasets
To perform the aforementioned tasks, large-scale document collections are essential for conducting self-supervised learning. Different pretraining datasets are adopted by various models. The most widely used pretraining dataset is the IIT-CDIP Test Collection 1.0 (Lewis et al., 2006), which contains more than 6 million documents with over 11 million scanned document images. Since it contains a cross-domain and large number of unannotated documents, it is used by the majority of models (Li et al., 2022b; Hong et al., 2021; Kim et al., 2022; Xu et al., 2020a, b; Huang et al., 2022; Xu et al., 2021; Bai et al., 2022; Appalaraju et al., 2021; Liao et al., 2023; Wang et al., 2022b; Gu et al., 2021; Zhai et al., 2023; Li et al., 2021a; Luo et al., 2023; Chen et al., 2022; Appalaraju et al., 2024). As the original IIT-CDIP dataset provides the text content without layout information, off-the-shelf OCR tools are normally used to acquire the bounding box information of each document. Some models (Li et al., 2021c, b; Wang et al., 2022a) use relatively smaller pretraining datasets like RVL-CDIP (Harley et al., 2015), which contains 400,000 evenly distributed documents in 16 types, to reduce the cost of pretraining. The multi-source model XDoc (Chen et al., 2022) also leverages many plain text corpora for pretraining, such as BookCORPUS, CC-NEWS, OPENWEBTEXT, STORIES and HTML-sourced CommonCrawl datasets. To address multilingual scenarios, both LiLT (Wang et al., 2022b) and LayoutXLM (Xu et al., 2021) follow the principles and policies of Common Crawl to gather large amounts of multilingual digitally-born PDF documents.
7. Conclusion
This paper comprehensively reviews deep learning-based models for visually rich document content understanding, encompassing both mono-task frameworks designed for specific VRDU downstream tasks and multi-task frameworks that support multiple VRDU downstream tasks. Beyond introducing the novelties of each model, the limitations of these frameworks are summarized at the end of each section, offering a thorough trend analysis. Additionally, this paper summarizes existing VRD content understanding datasets, pointing out future trends and demands for VRDU. To provide a systematic review, we critically discuss various techniques, highlighting their strengths and limitations. We believe that this survey offers a comprehensive overview of the development of visually rich document content understanding, catering to the needs of both the academic and industrial sectors.

### v2

@@ 6.4.2 Pretraining Datasets
To perform the aforementioned tasks, large-scale document collections are essential for conducting self-supervised learning. Different pretraining datasets are adopted by various models. The most widely used pretraining dataset is the IIT-CDIP Test Collection 1.0 [iitcdip], which contains more than 6 million documents with over 11 million scanned document images. Since it contains a cross-domain and large number of unannotated documents, it is used by the majority of models [dit, bros, donut, layoutlm, layoutlmv2, layoutlmv3, layoutxlm, wukong, docformer, doctr, lilt, udoc, faststructext, structurallm, geolayoutlm, xdoc, docformerv2]. As the original IIT-CDIP dataset provides the text content without layout information, off-the-shelf OCR tools are normally used to acquire the bounding box information of each document. Some models [structext, selfdoc, mgdoc] use relatively smaller pretraining datasets like RVL-CDIP [rvlcdip], which contains 400,000 evenly distributed documents in 16 types, to reduce the cost of pretraining. The multi-source model XDoc [xdoc] also leverages many plain text corpora for pretraining, such as BookCORPUS, CC-NEWS, OPENWEBTEXT, STORIES and HTML-sourced CommonCrawl datasets. To address multilingual scenarios, both LiLT [lilt] and LayoutXLM [layoutxlm] follow the principles and policies of Common Crawl to gather large amounts of multilingual digitally-born PDF documents.
6.5 LLM and MLLM into VRD-CU
In addition to frameworks specifically tailored for VRD-CU tasks, a growing number of general-purpose MLLMs have been proposed. These models aim to handle a wide range of modalities and domains—including plain text, natural scenes, documents, charts, and tables—and have demonstrated strong performance on various VRD-CU benchmarks. Many of these models incorporate specialized architectural mechanisms, pretraining or instruction tuning to support broad vision-language understanding across tasks. These general-purpose models can be broadly classified into two categories based on their reliance on text extraction: OCR-dependent and OCR-free LLMs/MLLMs.
6.5.1 OCR-Dependent LLMs/MLLMs
OCR-dependent LLMs/MLLMs rely on off-the-shelf OCR tools to extract text and layout information, typically in the form of tokens and bounding boxes, from scanned or digital documents. These extracted elements are either directly embedded into LLM prompts (e.g., ICL-D3IE [icld3ie], DocLLM [docllm]) or fused with visual features using multimodal encoders like ViT [vit] or LayoutLMv3 [layoutlmv3] (e.g., LayoutLLM [layoutllm], DoCo [li2024enhancing], GPE [gpe]). To incorporate layout cues, models commonly use 2D positional encodings, bounding box-aware prompts, or auxiliary encoders, with some frameworks verbalizing spatial coordinates or using layout-specific tokens to enrich input representations. While this modular design reduces the need for costly end-to-end pretraining on text recognition, it introduces vulnerabilities: OCR errors, especially in noisy or handwritten documents, can cascade through the pipeline, and unordered extracted text often requires additional processing like XY-cut reordering. Furthermore, layout representations are often implicit or rule-based [lapdoc, icld3ie], which may fail to capture fine-grained structural dependencies. The use of low-resolution images to reduce computation can also limit the expressiveness of visual features, leading to degraded multimodal fusion and overall performance, especially on layout-sensitive tasks.
6.5.2 OCR-Free LLMs/MLLMs
OCR-free LLMs/MLLMs have emerged as a promising direction when applying in end-to-end VRD-CU, eliminating the dependency on explicit text extraction by operating directly on document images. These models typically employ one or more vision encoders to extract visual features, which are then fused with user queries and decoded by LLMs to generate task-specific outputs such as, Donut, mPLUG-DocOwl, and UReader [donut, mplugdocowl, ureader]. Achieving fine-grained textual understanding in OCR-free settings requires high-resolution input, which results in long visual sequences and necessitates the use of visual compression modules to maintain computational tractability [hrvda, mplugdocowl15]. Additionally, effective integration of textual semantics and document layout often hinges on large-scale pretraining. This is typically achieved through auxiliary tasks such as text recognition [textmonkey] and image captioning [feng2024docpedia], which facilitate multimodal alignment. However, the paradigm is heavily reliant on extensive dataset construction and significant computational resources, presenting considerable challenges for scalability and real-world deployment.
6.6 Performance Analysis